# Benchmarking AI Models in Software Engineering: A Review, Search Tool, and Unified Approach for Elevating Benchmark Quality

### v2

@@ 6.5. Threats to the Validity
External Validity: A threat here is the generalizability of the user study’s results. We mitigated this by including 22 participants, but the sample size may still limit the generalizability of the results. Although we evaluated the effects of the BenchFrame framework on the performance of ten models with one of the most widely used benchmarks, this could not be considered sufficient to prove the generalizability of the results; adding more models and application to more benchmarks could further confirm our results. BenchFrame may be criticized for its limited applicability due to the significant manual effort it demands compared to more automated systems like EvalPlus. We demonstrate that the process of improving benchmarks such as HumanEval and MBPP can be (semi-)automated through the use of agentic pipelines.
7. Related Work
@@
Initial research in this area included broad surveys of LLMs in software engineering, which established evaluation as a central research challenge (Fan et al., 2023b; Hou et al., 2024; Zakeri-Nasrabadi et al., 2023; Zhang et al., 2023c). More recently, meta-analyses have focused on the benchmarks themselves. For instance, Wang et al. created a taxonomy based on the Software Development Life Cycle, revealing a research gap with a strong focus on code implementation and a negligence towards design and requirements engineering (Wang et al., 2025a). Our work contributes to this area by providing a comprehensive review of 273 benchmarks from 247 studies (RQ1). We go beyond categorization by analyzing each benchmark across a 14-point metadata schema to identify systemic limitations, such as poor maintenance, language specificity, and a lack of peer review.
@@
While platforms like Hugging Face and the now-discontinued Papers With Code serve as valuable repositories for hosting datasets (Hugging Face, 2016), they often lack the specialized search and visualization capabilities needed by AI4SE researchers. BenchScout addresses this gap (RQ2) by providing an extensible semantic search tool specifically designed for the AI4SE community. Unlike general-purpose repositories, it combines semantic embeddings with structured metadata filters and an interactive visualization of the benchmark landscape to improve the discoverability and selection of relevant evaluation tools.
@@
There is a growing agreement that many static benchmarks are prone to saturation and data contamination, which can compromise the validity of evaluation results (Jain et al., 2024). In response, two movements have arisen. Efforts like LiveCodeBench fall under the ”build anew” category; these strive to develop dynamic, ”live” benchmarks that continuously gather new challenges from real-world environments to avoid overfitting.
Our method, BenchFrame, belongs to the ”repair and refine” category and focuses on improving the quality of existing benchmarks (RQ3). A method within this category is DyCodeEval, which uses seed contexts to frame problems in specific scenarios (e.g., banking, healthcare, education). This approach is designed to produce more reliable results by minimizing data contamination and memorization effects (Chen et al., 2025). Our approach sets itself apart by incorporating a peer-review-oriented process that includes standardized observations, precise modifications, and independent validation. Like DyCodeEval, we show its practical scalability through an agentic pipeline that automates the improvement process and tackles the main challenge, namely, the manual efforts required in BenchFrame.
8. Conclusion

### v3

@@ VI-E Threats to the Validity
External Validity: A threat here is the generalizability of the user study’s results. We mitigated this by including 22 participants, but the sample size may still limit the generalizability of the results. Although we evaluated the effects of the BenchFrame approach on the performance of ten models with one of the most widely used benchmarks, this could not be considered sufficient to prove the generalizability of the results; adding more models and application to more benchmarks could further confirm our results. BenchFrame may be criticized for its limited applicability due to the significant manual effort it demands compared to more automated systems like EvalPlus. We demonstrate that the process of improving benchmarks such as HumanEval and MBPP can be (semi-)automated through the use of agentic pipelines.
VII Related Work
@@
Initial research in this area included broad surveys of LLMs in software engineering, which established evaluation as a central research challenge [248, 249, 250, 174]. More recently, meta-analyses have focused on the benchmarks themselves. For instance, Wang et al. created a taxonomy based on the Software Development Life Cycle, revealing a research gap with a strong focus on code implementation and a negligence towards design and requirements engineering [251]. Our work contributes to this area by providing a comprehensive review of 273 benchmarks from 247 studies (RQ1). We go beyond categorization by analyzing each benchmark across a 14-point metadata schema to identify systemic limitations, such as poor maintenance, language specificity, and a lack of peer review.
@@
While platforms like Hugging Face and the now-discontinued Papers With Code serve as valuable repositories for hosting datasets [252], they often lack the specialized search and visualization capabilities needed by AI4SE researchers. BenchScout addresses this gap (RQ2) by providing an extensible semantic search tool specifically designed for the AI4SE community. Unlike general-purpose repositories, it combines semantic embeddings with structured metadata filters and an interactive visualization of the benchmark landscape to improve the discoverability and selection of relevant evaluation tools.
@@
There is a growing agreement that many static benchmarks are prone to saturation and data contamination, which can compromise the validity of evaluation results [18]. In response, two movements have arisen. Efforts like LiveCodeBench fall under the ”build anew” category; these strive to develop dynamic, ”live” benchmarks that continuously gather new challenges from real-world environments to avoid overfitting.
Our method, BenchFrame, belongs to the ”repair and refine” category and focuses on improving the quality of existing benchmarks (RQ3). A method within this category is DyCodeEval, which uses seed contexts to frame problems in specific scenarios (e.g., banking, healthcare, education). This approach is designed to produce more reliable results by minimizing data contamination and memorization effects [253]. Our approach sets itself apart by incorporating a peer-review-oriented process that includes standardized observations, precise modifications, and independent validation. Like DyCodeEval, we show its practical scalability through an agentic pipeline that automates the improvement process and tackles the main challenge, namely, the manual efforts required in BenchFrame.
VIII Conclusion
@@ VIII-A Data Availability
IX Acknowledgments
This research was supported in part by an Amazon Research Award granted to Dr. Maliheh Izadi. We gratefully acknowledge Amazon’s support. The views and conclusions contained in this paper are those of the authors and do not necessarily reflect the position or policies of Amazon.